OpenStudy (anonymous):

A rifle is aimed horizontally at a target 50 m away. The bullet hits the target 2.6 cm below the aim point. (a) What was the bullet's flight time? s (b) What was the bullet's speed as it left the barrel? m/s

OpenStudy (souvik):

can u say why the bullet hits bellow the target?

OpenStudy (kainui):

So in order to solve this problem you need to consider the equations of motion which look like this: \[s_f=s_i+v_i t+\frac{1}{2}at^2\] So here I have: s=distance v=velocity a=acceleration t=time We have movement in two directions, the x and y directions, so we have two equations! \[x_f=x_i+v_{ix} t+\frac{1}{2}a_{x}t^2\]\[y_f=y_i+v_{iy} t+\frac{1}{2}a_yt^2\] So now let's draw a picture and label things and see if we can figure out what the acceleration, velocity, distance, and time is.  Certain things go away, like the initial y-velocity is 0 right? Also there is no acceleration in the x-direction!